e2:

Which linear function represents the line given by the point-slope equation y + 1 = –3(x – 5)?

e2:

does anyone know this problem

Nnesha:

\[\rm y-y_1=m(x-x_1)\] is the general form, where `m` is the slope and (x_1, y_1) represents the coordinate of one of the points on the line.

Nnesha:

so what is the slope in this equation ?\[y + 1 = –3(x – 5)\]

e2:

idk tbh i am barely going back to school

e2:

i am a teen mom

Nnesha:

what dod you mean by this ` i am barely going back to school`? if you are a teen mom then it is absolutely necessary to get a degree not only for yourself but also for a better future for your kids

Nnesha:

\[\rm y\color{orange}{-y_1}=\color{red}{m}(x\color{blue}{-x_1})\] \[y \color{orange}{ +1} =\color{red}{ –3}(x ~\color{blue}{–~ 5})\] m represents the slope. m is replaced by what number in 2nd equation ?